OpenStudy (anonymous):

Prove the Pythagorean Theorem using similar triangles. The Pythagorean Theorem states that in a right triangle, the sum of the squares of the lengths of the legs of the triangle equals the squared length of the hypotenuse. Be sure to create and name the appropriate geometric figures.

OpenStudy (anonymous):

by AA similarity postulate: △ADB∼△BDC∼△ABC therefore the sides of the triangles are proportional: AD/AB = AB/AC BC/AC = DC/BC By cross multiplication we have the following equations: AD * AC = AB * AB AC * DC = BC * BC Which is the same as: AD * DC = BD^2 AD * AC = AB^2 AC * DC = BD^2 Add the two equations: AD * AC AC * DC = AB^2 BC^2 By the distributive property: AC(AD DC) = AB^2 BC^2 but by construction AD DC = AC...so we have: AC * AC = AB^2 BC^2 Which is equivalent to: AB^2 BC^2 = AC^2 Which is what we want to prove.

OpenStudy (anonymous):

that is what i have... i just want to make sure im right..

OpenStudy (anonymous):

i love this i got it

OpenStudy (anonymous):

well a^2+b^2=c^2 so is that what you have as your first step

OpenStudy (anonymous):

im guessing..the coding got messed up and idk what it is now XD

OpenStudy (anonymous):

you used the distubitive propery i think

OpenStudy (anonymous):

So do u think that is a correct proof to answer the question?

OpenStudy (anonymous):

i'm guessing, probably

OpenStudy (anonymous):

no do u know how to do the theroum

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

awwww ship

OpenStudy (anonymous):

wanna learn

OpenStudy (anonymous):

...i guess?

OpenStudy (anonymous):

i just wanted to know if what i did was right or not..

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

well not u did the distributive property the Pythagorean Theorem is a^2+b^2=c^2 ill give u the site to understand it better

OpenStudy (anonymous):

https://en.wikipedia.org/wiki/Pythagorean_theorem go to this site

OpenStudy (anonymous):

So should i just change the top line to the "a^2+b^2=c^"??

OpenStudy (anonymous):

then your answer will be all wrong

OpenStudy (anonymous):

0-0 okay.. then help me re-do it?

OpenStudy (anonymous):

ight

OpenStudy (anonymous):

Awesome :D

OpenStudy (anonymous):

@teagirl7630

OpenStudy (anonymous):

so as you know a^2+b^2=C^2

OpenStudy (anonymous):

okay, so that's how we begin?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

What's next?

OpenStudy (anonymous):

it A

OpenStudy (anonymous):

then on your tiangle it shows all the and god for flam it love2hot2 and it will say what A, B or c is

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

...huh?

OpenStudy (anonymous):

gtg ill message you

OpenStudy (anonymous):

.-. mmkay

OpenStudy (anonymous):

yeah

OpenStudy (ybarrap):

I'm doing this because I'm not sure how to read what you wrote, so we are starting from scratch. And if you know what to do next. Just do it.

OpenStudy (anonymous):

not really.. i think i just need the two colomn proof..but im not sure how to do it correctly.

OpenStudy (ybarrap):

Follow my steps here in the order I show, referring to the diagram.  We use similarity of 3 triangles to prove. Does this make sense?